OpenStudy (anonymous):

-x+3y=6 what would the graph look like on this matter and then what would be x if y=1

OpenStudy (anonymous):

Why don't you make a table for x and y so you can find the coordinates and then plot them on the graph

OpenStudy (anonymous):

I am trying to figure that out but it seems like I am doing everything wrong

OpenStudy (anonymous):

OK lets take a range from -3 to 3

OpenStudy (nubeer):

well if u had made the table.. please post it here so we all can see where u went wrong.

OpenStudy (anonymous):

Re arrange the formula so you can be capable of graphing it. The resulting arrangement would be : y=(x/3)+2 Vertical translation two units up and horizontal stretch by a factor of two. It also has y-intercept at +2 As if you plug in 1 for y, you would get -3 for x. (-3,1)

OpenStudy (anonymous):

nubeer that was my issue I could not even do the problem, I am so confused on it

OpenStudy (anonymous):

thanks boblovesmath, I wish that I loved math because I have been trying to do this final for 7 hours and only half way done

OpenStudy (anonymous):

 Just complete the table

OpenStudy (nubeer):

Faris has made the table, you have the equation y=(x/3) +2 start finding the value of y for each value of x in the table.. complete the table and plot all the points in table on the graph.. that will do it for you.

OpenStudy (nubeer):

for example.. if we take x=-3 from the table drawn the value of y will be y =(-3/3)+2 y = -1+2 y=1 so u just put 1 in table under value of y same way u have to find for the rest

OpenStudy (nubeer):

@bigmommatiad